OpenStudy (anonymous):

so s =4b ----------------------> equation 1 s+6 = 3(b+6) ----------------> equation 2 solve the two equations and find s and b

OpenStudy (anonymous):

substitute the value of s=4b in the equation s+6=3(b+6) 4b+6=3b+18 4b-3b=18-6 b=12 s=4b s=4(12) s=48

OpenStudy (anonymous):

i hope i have helped you.^_^